OpenStudy (anonymous):

a right triangle has a leg measuring 52m and hypotenuse 63m. find the other leg? pleaseeeeee

OpenStudy (anonymous):

use the pythagorean theorem

OpenStudy (anonymous):

do u know how do it. can you do it for me step by step

OpenStudy (anonymous):

other leg = \[\sqrt{62^{2}-52^{2}}\]

OpenStudy (anonymous):

Solve.

OpenStudy (anonymous):

see... in right triangle one leg^2 + second leg^2 = hypotenuse^2

OpenStudy (anonymous):

63*

OpenStudy (anonymous):

u know one leg and the hypotenuse...plug those in and find thr second leg...

OpenStudy (anonymous):

\[\sqrt{63^{2}-52^{2}}\]

OpenStudy (anonymous):

i got 35.56683849

OpenStudy (anonymous):

\[\sqrt{63^{2} - 52^{2}}\]

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

round it it would be 36